OpenStudy (anonymous):

find the general solution of the differential equation, xy'+y=1+x.

OpenStudy (kainui):

Can you find the solution to the homogeneous version?

OpenStudy (anonymous):

If i can find any I wouldn't ask

OpenStudy (kainui):

Good luck then.

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

use d(xy) =xdy+ydx

OpenStudy (anonymous):

but x is separate, it's like x·y'+y=1+x

OpenStudy (anonymous):

xy'+y=1+x can be stated as xdy/dx +y = 1+x ==>> xdy +ydx =(1+x)dx d(xy) = (1+x)dx

OpenStudy (anonymous):

now integrate both sides.. we have xy = x+x^2 /2 +c

OpenStudy (anonymous):

oh i got it thx!!

OpenStudy (anonymous):

welcome dear....